The train stations for New York and Boston are 220 miles apart. A train left New York for Boston at 8:00 am traveling at 65 mph. An hour later an express train left Boston for New YOrk traveling at 90 mph. At what time did they pass each other?


Answer by mananth(16949)

Train A New York Boston 8.00 am- 65 mph -
Train B Bost New York 9.00 am - 90 mph
..
Train A in 1 hour before train B started has traveled 65 miles.
Balnce distance = 220 -65 = 155 miles.
..
At this point distance = 155 miles.
travelling towards each other.
So add the speeds 90+65= 155 mph
..
t = d/r
t = 155/155 mph
..
t = 1hour after Train B leaves.
So at 10.00 am they cross each other.
..
m.ananth@hotmail.ca